An investor has an investment capital of KES. 9,000,000. He wishes to invest in a portfolio of two securities, A and B in the following proportion; KES. 4,950,000 in security A and KES. 4,050,000 in security B.

The returns on these two securities depend on the state of the economy as shown below:

State of Economy

Probability

Return on Security A

Return on security B

Expansion

0.4

25%

5%

Boom

0.3

20%

10%

Recession

0.2

12%

20%

Depression

0.1

-20%

35%

Required:

  1. Compute the expected portfolio return.
  2. Determine the correlation coefficient between security A and Security B and interpret it.
  3. Calculate the portfolio risk as measured by standard deviation.
Answer #1
Total Portfolio value = Value of Sec A + Value of Sec B
=4950000+4050000
=9000000
Weight of Sec A = Value of Sec A/Total Portfolio Value
= 4950000/9000000
=0.55
Weight of Sec B = Value of Sec B/Total Portfolio Value
= 4050000/9000000
=0.45

i

Sec A
Scenario Probability Return% =rate of return% * probability Actual return -expected return(A)% (A)^2* probability
Expansion 0.4 25 10 8.6 0.0029584
Boom 0.3 20 6 3.6 0.0003888
Recession 0.2 12 2.4 -4.4 0.0003872
Depression 0.1 -20 -2 -36.4 0.0132496
Expected return %= sum of weighted return = 16.4 Sum=Variance Sec A= 0.01698
Standard deviation of Sec A% =(Variance)^(1/2) 13.03
Sec B
Scenario Probability Return% =rate of return% * probability Actual return -expected return(A)% (B)^2* probability
Expansion 0.4 5 2 -7.5 0.00225
Boom 0.3 10 3 -2.5 0.0001875
Recession 0.2 20 4 7.5 0.001125
Depression 0.1 35 3.5 22.5 0.0050625
Expected return %= sum of weighted return = 12.5 Sum=Variance Sec B= 0.00863
Standard deviation of Sec B% =(Variance)^(1/2) 9.29
Expected return%= Wt Sec A*Return Sec A+Wt Sec B*Return Sec B
Expected return%= 0.55*16.4+0.45*12.5
Expected return%= 14.65

ii

Covariance Sec A Sec B:
Scenario Probability Actual return% -expected return% for A(A) Actual return% -expected return% For B(B) (A)*(B)*probability
Expansion 0.4 8.6 -7.5 -0.00258
Boom 0.3 3.6 -2.5 -0.00027
Recession 0.2 -4.4 7.5 -0.00066
Depression 0.1 -36.4 22.5 -0.00819
Covariance=sum= -0.0117
Correlation A&B= Covariance/(std devA*std devB)= -0.97

iii

Variance =( w2A*σ2(RA) + w2B*σ2(RB) + 2*(wA)*(wB)*Cor(RA, RB)*σ(RA)*σ(RB))
Variance =0.55^2*0.13032^2+0.45^2*0.09287^2+2*0.55*0.45*0.13032*0.09287*-0.96669
Variance 0.00109
Standard deviation= (variance)^0.5
Standard deviation= 3.30%